# Efficient and Transferable Agentic Knowledge Graph RAG via Reinforcement Learning

### v3

@@ 4.2 Cross-KG Transfer via Plug-and-Play
Table 2: Zero-shot cross-KG transferability of KG-R1. Agents are trained on WebQSP or CWQ and evaluated on new benchmarks by swapping only the KG-backend server (no policy retraining). AVG. is averaged across QA samples. The max turn number set to $H=5$. All scores are percentages.
Freebase-based Wikidata-based Temporal
Method Trained KG SimpleQA GrailQA T-REx QALD-10en MultiTQ AVG.
F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1
LLM-only Methods
Vanilla Qwen 2.5 3B-IT 13.7 / 13.7 15.9 / 15.9 24.4 / 24.4 23.8 / 23.8 2.2 / 5.4 19.4 / 19.8
CoT Qwen 2.5 3B-IT 10.9 / 10.9 18.2 / 18.2 22.0 / 22.0 25.9 / 25.9 1.9 / 3.9 18.0 / 18.2
LLM+KG Methods
KG-specific Baselines
$79.6^{1}/80.6^{1}$
$76.1^{2}/77.1^{2}$
$84.4^{3}/79.1^{3}$
$64.9^{4}/59.6^{4}$
— / $85.1^{5}$
$49.8^{6}/50.8^{6}$
$62.2^{7}/63.2^{7}$
— / $79.7^{8}$
— / $38.0^{9}$
RoG 13.5 / 43.5 6.4 / 23.7 8.1 / 36.4 11.4 / 44.3 3.6 / 8.7 8.1 / 34
ToG 50.2 / 53.6 63.9 / 68.7 79.3 / 76.4 48.7 / 50.2 25.1 / 27.9 66.2 / 65.9
ReKnoS 50.7 / 52.9 65.8 / 71.9 81.4 / 78.9 50.6 / 53.8 28.5 / 31.1 68.2 / 68.2
KG-R1 (Our Methods)
KG-R1 (1 run) WebQSP 59.1 / 59.1 32.8 / 38.5 80.5 / 84.5 51.9 / 53.4 21.6 / 31.4 64.0 / 68.3
KG-R1 (1 run) CWQ 64.6 / 64.7 42.8 / 50.2 81.3 / 85.6 55.9 / 57.7 27.1 / 38.9 67.2 / 72.1
KG-R1 (3 runs) WebQSP 69.1 / 69.1 44.6 / 52.1 84.6 / 88.8 61.3 / 62.8 33.5 / 45.7 70.9 / 75.8
KG-R1 (3 runs) CWQ 73.1 / 73.1 52.8 / 61.0 86.8 / 91.5 63.9 / 65.5 36.2 / 48.4 74.1 / 79.4
1SPARKLE 2GETT-QA 3SG-KBQA 4ARG-KBQA 5ATLAS 6COT-SPARQL 7DFSL-MQ 8Prog-TQA 9ARI ; see Table 7 for references.
A4. KG-R1 transfers across diverse types of KGs. Table 2 reports zero-shot, plug-and-play results. For comparability, we report AVG., the F1/Hits@1 averaged across the five KGQA datasets, weighted by QA counts. KG-R1 (1 run) substantially outperforms LLM-only baselines using the same base model—Vanilla (19.4/19.8 F1/Hits@1) and CoT (18.0/18.2)—reaching 64.0/68.3 when trained on WebQSP and 67.2/72.1 when trained on CWQ. The gains hold not only on Freebase-based QA but also on Wikidata and Temporal benchmarks, indicating that reinforcement-learned KG-RAG transfer across diverse KGs rather than narrow, in-domain improvements.
A5. KG-R1 achieves accurate inference comparable to LLM+KG baselines. Against LLM+KG baselines, KG-R1 (N=1) is slightly better on average: ToG (66.2/65.9) and ReKnoS (68.2/68.2). Notably, KG-R1, with a 3B model, delivers higher Hits@1 than both ToG and ReKnoS that uses strong foundation models. Increasing to N=3 runs boosts performance to 74.1/70.4 (CWQ-trained), making KG-R1 comparable to KG-specific baselines on average. We observe that RoG collapses (0/0 average) revealing brittle, schema-specific planning with fine-tuned model that does not transfer. Overall, KG-R1 attains strong transferability among plug-and-play methods with a lightweight model, supporting its practicality for real-world KG-RAG deployment across diverse KGs.
A6. Training data on transferability. KG-R1 models trained on CWQ consistently outperform WebQSP-trained variants across all transfer targets, averaging 1–5% higher F1 and Hits@1. Given that CWQ provides a larger training set with greater multi-hop complexity than WebQSP, this suggests that exposure to more complex and diverse reasoning patterns yields superior generalization.
@@ 4.3 Ablation Studies
We conduct ablation studies for the three core components of KG-R1: reward design (Section 3.3), RL algorithm (Section 2), and KG-server implementation (Section 3) and base LLM size. We train Qwen-2.5-3B-it on WebQSP and CWQ with maximum turn number $H=5$ and report final step performance(F1/Hit@1/Ret.rate in Table 3). The training curves are shown in Figures 8, 9.
Table 3: Ablation studies of KG-R1 components on CWQ and WebQSP datasets.
Method WebQSP CWQ F1 Hit@1 Ret.Rate F1 Hit@1 Ret.Rate KG-R1 Default KG-R1 77.2 84.3 76.8 66.8 69.7 51.4 Reward Ablation w/o Turn Rewards 67.1 (-13.1%) 76.0 (-9.8%) 68.1 51.9 (-22.3%) 52.9 (-24.1%) 44.7 w/o Turnwise Advantage 49.6 (-35.8%) 61.2 (-27.4%) 22.8 63.5 (-4.9%) 64.5 (-7.5%) 49.4 w/o Retrieval Reward 45.8 (-40.6%) 57.7 (-31.5%) 18.6 66.0 (-1.2%) 67.0 (-3.9%) 47.2 Other Ablations w PPO 0.0 (-100.0%) 0.0 (-100.0%) 0.0 0.0 (-100.0%) 0.0 (-100.0%) 0.0 w/o Hierarchical Rel. Retr. 76.8 (-0.6%) 83.1 (-1.5%) 76.0 55.5 (-16.9%) 58.6 (-15.9%) 39.7 w Qwen-2.5-7B-it 79.6 (+3.1%) 86.6 (+2.7%) 83.4 (+8.6%) 65.6 (-1.8%) 68.7 (-1.4%) 58.8 (+14.4%)
PPO training crashed.    Peak values before 7B model training collapse.
A7. Reward function. First, removing turn-level rewards yields the largest drop on both WebQSP (13.1% / 9.9% in F1 score and Hit@1, respectively) and CWQ (22.3% / 24.1%), highlighting the importance of per-turn validity/effectiveness signals. Next, disabling the turn-wise advantage calculation produces a substantial decline on WebQSP (35.8% / 27.4%) but only a moderate drop on CWQ (4.9% / 7.5%), indicating that turn-specific learning signals matter more for certain datasets. Finally, removing the retrieval reward significantly impacts WebQSP (40.7% / 31.5%) but marginally affects CWQ (1.2% / 3.9%), though it substantially reduces the retrieval success rate of gold entities (Ret.Rate) on both datasets, suggesting that it is essential for motivating exploration of the KG.
A8. RL algorithm (GRPO vs. PPO). Replacing GRPO with vanilla PPO (Ouyang et al., 2022) results in learning collapse (0.0 across all metrics). Under PPO, we observe reward hacking (see examples in Appendix E.3.3): the agent fabricates ”retrieved” content matching expected formats to earn reward. Since PPO’s value critic (another LLM) cannot distinguish genuine KG retrievals from hallucinations, its advantage estimates become unstable, driving training collapse. This underscores the importance of relative advantage methods like GRPO for multi-step reasoning.
A9. Retrieval format. In our standard KG-R1 setup, one-hop relations retrieved by $\mathtt{get\_head\_relations}$ and $\mathtt{get\_tail\_relations}$ are appended to the agent’s context as a flat, comma-separated list (e.g., “rel1, rel2, …”). This can lead to excessive token consumption when many relations are retrieved. Because Freebase relations use a compositional format (e.g., “domain.type.property”), we tested a hierarchical relation retrieval that groups relations into a tree to reduce redundancy (see Appendix E.3.2). Our ablation studies show that switching from flat lists to a hierarchical format reduces performance—WebQSP shows 0.5% drop, while CWQ experiences 16.9% drop. These results suggest that despite its higher token usage, the flat retrieval format provides more direct access to relation information, which proves critical for training.
A10. Model scale. The 7B model improves more quickly than the 3B model but exhibits training collapse (around 350 steps on WebQSP and 200 steps on CWQ). This aligns with reports that fast distribution shift during GRPO—amplified by larger models and higher update rates—drives entropy collapse and model divergence (Jin et al., 2025; Liu et al., 2025).
@@ 5 Conclusion
We propose KG-R1, a single-agent, multi-turn KG-RAG framework in which one LLM queries a lightweight KG server and is optimized via RL. Across KG-augmented QA benchmarks, KG-R1 attains strong accuracy while using markedly fewer tokens and lower inference cost than prior work. It also supports plug-and-play transfer, retaining robust performance across KGs. Together, these results position KG-R1 as a practical and deployable KG-RAG system for real-world use.
Appendix
Appendix A Theoretical Proofs
A.1 Completeness and Transferability of the KG-Server
Here we provide robust proofs for completeness and transferability of $\mathcal{U}_{ret}$ in our KG retrieval server.
Preliminaries. Let $G=\{(e,r,e^{\prime})\mid e,e^{\prime}\!\in\!\mathcal{E},\,r\!\in\!\mathcal{R}\}$ be a directed KG. Define the KG-R1 action set
@@ A.1 Completeness and Transferability of the KG-Server
Proposition A.1 (Finite-Horizon Bound).
For any path $Z$ of length $\ell$, there exists an action sequence from $\mathcal{U}_{\text{ret}}$ of length at most $\ell$ whose final output contains $e_{\ell}$.
Proof. Along the path $Z$, for each step $i=1,\ldots,\ell$ the call $\mathtt{get\_tail\_entities}(e_{i-1},r_{i})$ returns a set that contains $e_{i}$. Therefore, after the $\ell$-th step the output contains $e_{\ell}$.
Proposition A.2 (Completeness (Integrity)).
Every finite reasoning path in $G$ can be realized by a finite sequence of actions from $\mathcal{U}_{\text{ret}}$ whose outputs include its terminal entity.
@@
Proof. Starting with $\ell=1$, if $(e_{0},r_{1},e_{1})\in G$ then $e_{1}\in\mathtt{get\_tail\_entities}(e_{0},r_{1})$. For the inductive step, assume the claim holds for length $\ell-1$. By the hypothesis we reach $e_{\ell-1}$; applying $\mathtt{get\_tail\_entities}(e_{\ell-1},r_{\ell})$ then returns a set containing $e_{\ell}$.
Proposition A.3 (Schema-Free Transferability).
The semantics of each $a\in\mathcal{U}_{\text{ret}}$ depend only on membership of triples $(e,r,e^{\prime})$ in $G$. Replacing $G$ with any directed KG $G^{\prime}$ preserves action meaning and allows a fixed client policy to transfer unchanged.
Proof. Each operator is defined as a set comprehension over triples in $G$, independent of schema, datatypes, or ontology rules. Under $G\mapsto G^{\prime}$, the same comprehensions define the actions on $G^{\prime}$, so a learned policy remains well-defined.
Proposition A.4 (Minimality of $\mathcal{U}_{\text{ret}}$).
No proper subset of $\mathcal{U}_{\text{ret}}$ supports both symmetric navigation and local relation discovery; any operator that only filters or composes these four is derivable.
Proof. Removing $\mathtt{get\_tail\_entities}$ (resp. $\mathtt{get\_head\_entities}$) prevents forward (resp. backward) traversal. Removing both $\mathtt{get\_tail\_relations}$ and $\mathtt{get\_head\_relations}$ blocks local relation discovery when the relation set is unknown. Conversely, a conditional/filtered variant can be constructed by enumerating $\mathtt{get\_tail\_relations}(e)$ and intersecting results of $\mathtt{get\_tail\_entities}$; hence no additional primitives are required.
Appendix B Datasets
B.1 KGQA Datasets
These are the KGQA datasets we used in our work.
Freebase
@@ B.1 KGQA Datasets
WebQSP (Yih et al., 2016) (full test; 1,639) Open-domain questions with mostly 1–2 hop reasoning over Freebase MIDs and SPARQL annotations.
@@
ComplexWebQuestions (CWQ) (Talmor & Berant, 2018) (full test; 3,531) Compositional multi-hop questions generated from SPARQL templates that stress longer chains and constraint composition; used to probe multi-turn retrieval quality and robustness without dataset-specific rules.
@@
SimpleQuestions (SimpleQA) (Bordes et al., 2015) (1,000 test) Single-relation 1-hop questions over Freebase; serves as a retrieval-fidelity and token-efficiency baseline for KG-R1. We randomly sample 1,000 QA from the original test split (21,687).
@@
GrailQA (Gu et al., 2021) (1,000 test) Diverse compositional questions emphasizing generalization under Freebase; handled with the same minimal action interface and no hand-crafted schemas. We randomly sample 1,000 from the original test split (13,231).
Wikidata
@@
T-REx (Elsahar et al., 2018) (5,000 test) Large-scale slot-filling–style QA grounded in Wikidata triples; used to assess scalability and coverage under a different KG schema. We randomly sample 5,000 from the corpus ($\sim$2.2M).
@@
QALD-10en (Usbeck et al., 2023) (333 test) Manually curated, linguistically varied questions over Wikidata; useful for evaluating precision on a small but challenging set. We evaluate on 333 examples.
Temporal KG
MultiTQ (Chen et al., 2023) (1,000 test) Time-aware QA requiring temporal qualifiers (e.g., during, from–to); evaluates KG-R1 on temporal reasoning with time-scoped entities and relations. We randomly sample 1,000 from the original test split (9,000).
B.2 Dataset Preprocessing

### v5

@@ 4.4 RQ2: Cross-KG Transferability
Table 2: Zero-shot cross-KG transferability of KG-R1 (Hits@1, %). Agents are trained on WebQSP or CWQ and evaluated on new benchmarks by swapping only the KG-backend server (no policy retraining). AVG is averaged across QA samples. The maximum turn number is set to $H=5$.
Freebase-based Wikidata-based
[…]
Qwen 2.5-7B IO 10.30 $\pm$ 0.10 8.07 $\pm$ 0.06 22.79 $\pm$ 0.05 17.43 $\pm$ 0.01 18.84
CoT 7.07 $\pm$ 0.06 7.73 $\pm$ 0.12 17.38 $\pm$ 0.04 21.05 $\pm$ 0.06 14.82
SC 8.70 $\pm$ 0.20 8.60 $\pm$ 0.44 20.01 $\pm$ 0.12 20.32 $\pm$ 0.46 16.93
RoG 12.97 $\pm$ 0.35 17.67 $\pm$ 0.15 22.37 $\pm$ 0.27 29.33 $\pm$ 0.46 20.76
ToG 52.37 $\pm$ 0.80 54.73 $\pm$ 4.62 58.55 $\pm$ 1.11 50.95 $\pm$ 0.92 56.84
PoG 66.38 $\pm$ 0.39 69.88 $\pm$ 0.84 75.69 $\pm$ 0.08 46.07 $\pm$ 0.25 72.28
ReknoS 56.43 $\pm$ 1.22 69.58 $\pm$ 0.35 78.17 $\pm$ 0.03 37.94 $\pm$ 1.21 72.21
Qwen 2.5-72B IO 14.33 $\pm$ 0.15 11.63 $\pm$ 0.12 26.16 $\pm$ 0.02 17.45 $\pm$ 0.06 22.17
CoT 11.07 $\pm$ 0.12 10.53 $\pm$ 0.06 24.48 $\pm$ 0.02 20.52 $\pm$ 0.17 20.57
SC 11.60 $\pm$ 0.44 11.43 $\pm$ 0.25 26.01 $\pm$ 0.09 21.62 $\pm$ 0.60 21.86
RoG 20.40 $\pm$ 0.26 22.63 $\pm$ 4.51 32.50 $\pm$ 0.09 44.94 $\pm$ 0.69 30.07
ToG 57.83 $\pm$ 0.29 71.53 $\pm$ 0.76 67.93 $\pm$ 0.18 59.66 $\pm$ 0.46 66.67
PoG 65.67 $\pm$ 0.42 75.30 $\pm$ 0.99 79.52 $\pm$ 0.08 48.18 $\pm$ 3.53 75.63
ReknoS 68.40 $\pm$ 5.15 73.67 $\pm$ 0.60 77.11 $\pm$ 0.02 39.94 $\pm$ 0.30 73.77
Qwen 3-235B IO 11.37 $\pm$ 0.12 12.80 $\pm$ 0.10 34.61 $\pm$ 0.01 24.12 $\pm$ 0.17 27.99
CoT 14.40 $\pm$ 0.36 14.60 $\pm$ 0.30 37.72 $\pm$ 0.18 28.43 $\pm$ 0.17 30.97
SC 14.73 $\pm$ 0.42 15.73 $\pm$ 0.15 38.79 $\pm$ 0.27 29.53 $\pm$ 0.35 31.94
RoG 18.93 $\pm$ 0.51 18.10 $\pm$ 0.20 37.31 $\pm$ 0.19 37.54 $\pm$ 1.20 32.19
ToG 41.27 $\pm$ 3.61 53.20 $\pm$ 0.53 77.00 $\pm$ 0.10 57.76 $\pm$ 1.05 68.01
PoG 67.90 $\pm$ 0.17 77.93 $\pm$ 0.45 78.75 $\pm$ 0.13 52.55 $\pm$ 1.04 75.97
ReknoS 68.01 $\pm$ 3.37 75.67 $\pm$ 0.74 77.05 $\pm$ 0.03 39.44 $\pm$ 0.35 73.92
Qwen2.5 KG-R1 (WebQSP trained) 69.18 $\pm$ 0.10 68.55 $\pm$ 0.07 83.38 $\pm$ 0.08 41.96 $\pm$ 0.05 77.50
3B KG-R1 (CWQ trained) 57.31 $\pm$ 0.08 64.84 $\pm$ 0.04 80.10 $\pm$ 0.05 34.74 $\pm$ 0.12 72.82
Zero-shot transfer across unseen KGs. Table 2 reports zero-shot plug-and-play results, where the learned policy is directly applied to unseen KGs by swapping only the backend graph. For comparability, we report AVG, the Hits@1 averaged across all evaluation QA samples. KG-R1 demonstrates strong cross-KG transferability across both Freebase- and Wikidata-based benchmarks. When trained on WebQSP, KG-R1 achieves an average Hits@1 of 77.50, substantially outperforming LLM-only baselines (IO: 18.84, CoT: 14.82, SC: 16.93) and surpassing prior KG-augmented methods, including ToG (56.84), PoG (72.28), and ReKnoS (72.21). The gains are consistent across all transfer benchmarks, with particularly strong improvements on T-REx (83.38 vs. 78.17 for ReKnoS).
When trained on CWQ, KG-R1 achieves 72.82 Hits@1, remaining clearly above LLM-only baselines and competitive with strong KG-RAG methods built on much larger foundation models. In particular, the WebQSP-trained KG-R1 also outperforms ToG, PoG, and ReKnoS when these methods use Qwen3-235B backbones (77.50 vs. 68.01, 75.97, and 73.92, respectively), demonstrating that learned KG interaction policies can generalize better than simply scaling model size.
Notably, transfer gains are especially pronounced on Wikidata-based benchmarks (T-REx and QALD-10en), where schema differences are larger. This suggests that the schema-agnostic action design of KG-R1 enables robust generalization across diverse KG structures.
@@ 4.5 Ablation Studies
We ablate four key components of KG-R1: reward design, RL algorithm, retrieval format, and model scale. We train on WebQSP and CWQ with turn number $H=5$ and report final Hits@1 and retrieval success rate (Ret. Rate), defined as the fraction of queries where any ground-truth answer entity appears in the retrieved entities, in Table 3. The training curves are shown in Figure 7.
Table 3: Ablation studies of KG-R1 components.
Datasets $\rightarrow$ WebQSP CWQ
[…]
Model Scale. Scaling the base model from Qwen2.5-3B to 7B does not improve performance (WebQSP: 82.75 vs. 82.44; CWQ: 65.31 vs. 65.85), indicating that KG-R1’s gains mainly come from reinforcement-learned retrieval policies rather than model size alone.
@@ 5 Conclusion
We propose KG-R1, a single-agent, multi-turn KG-RAG framework where an LLM queries a lightweight KG server and is optimized via RL. Across KG-augmented QA benchmarks, KG-R1 achieves strong accuracy while using markedly fewer tokens and lower inference cost than in prior work. It also supports plug-and-play transfer, maintaining robust performance across KGs. Together, these results position KG-R1 as a practical and deployable KG-RAG system for real-world use.
Appendix A Discussion
Societal Impact.
KG-R1 aims to improve knowledge-grounded reasoning by enabling language models to interact with structured knowledge graphs through learned retrieval policies. This may benefit applications requiring factual consistency and traceable reasoning, such as question answering, scientific knowledge access, and enterprise knowledge systems. By reducing reliance on complex multi-module pipelines, KG-R1 may also improve the accessibility and deployment efficiency of knowledge-grounded reasoning systems.
Potential Negative Impact.
Like other retrieval-augmented systems, KG-R1 may propagate incorrect or outdated knowledge if the underlying knowledge graph is incomplete, noisy, or maliciously manipulated. In high-stakes domains such as medicine or finance, such errors could lead to misleading outputs. Additionally, improved knowledge-grounded reasoning may be misused to automate large-scale misinformation generation with higher factual coherence.
Limitations.
KG-R1 assumes access to a reasonably structured and queryable knowledge graph, which may not be available in all domains. Its performance depends on the quality and schema consistency of the KG, and transferability may degrade under severe schema mismatch or noisy graph structures. Furthermore, while KG-R1 improves inference efficiency through structured interaction, reinforcement learning training remains computationally expensive compared to standard supervised fine-tuning.
Appendix B Completeness and Transferability of the KG-Server
Here, we discuss the completeness and transferability of $\mathcal{U}_{ret}$ in our KG retrieval server.
Preliminaries.
Let $G=\{(e,r,e^{\prime})\mid e,e^{\prime}\!\in\!\mathcal{E},\,r\!\in\!\mathcal{R}\}$ be a directed KG. Define the KG-R1 action set
@@ Appendix B Completeness and Transferability of the KG-Server
[…]
@@
Proof. Starting with $\ell=1$, if $(e_{0},r_{1},e_{1})\in G$ then $e_{1}\in\mathtt{get\_tail\_entities}(e_{0},r_{1})$. For the inductive step, assume the claim holds for length $\ell-1$. By the induction hypothesis, we reach $e_{\ell-1}$; applying $\mathtt{get\_tail\_entities}(e_{\ell-1},r_{\ell})$ then returns a set containing $e_{\ell}$.
Schema-Free Transferability.
The retrieval action interface $\mathcal{U}_{\text{ret}}$ is defined purely in terms of set operations over triples $(e,r,e^{\prime})$, independent of any schema, ontology, or datatype constraints. As a result, the same action interface can be applied to different knowledge graphs without modification. This allows a learned policy to be executed on a new KG by simply replacing the backend graph.
However, this property guarantees only that the policy remains well-defined under KG replacement. The effectiveness of transfer depends on the similarity between the training and target KGs and is evaluated empirically in Section 4.4.
Appendix C Datasets
C.1 KGQA Datasets
Here, we list the KGQA datasets used in this paper.
Freebase-based Datasets
@@ C.1 KGQA Datasets
  • WebQSP [31] (1,639 test QA pairs) Open-domain questions with mostly 1–2 hop reasoning over Freebase MIDs and SPARQL annotations.
@@
  • ComplexWebQuestions (CWQ) [24] (3,531 test QA pairs) Compositional multi-hop questions generated from SPARQL templates that stress longer chains and constraint composition are used to probe multi-turn retrieval quality and robustness without dataset-specific rules.
@@
  • SimpleQuestions (SimpleQA) [2] (1,000 sampled test QA pairs) Single-relation 1-hop questions over Freebase serve as a retrieval-fidelity and token-efficiency baseline for KG-R1. We randomly sample 1,000 QA from the original test split (21,687).
@@
  • GrailQA [11] (1,000 sampled test QA pairs) Diverse compositional questions emphasizing generalization under Freebase; handled with the same minimal action interface and no hand-crafted schemas. We randomly sample 1,000 from the original test split (13,231).
Wikidata-based Datasets
@@
  • T-REx [9] (5,000 test QA pairs) Large-scale slot-filling–style QA grounded in Wikidata triples; used to assess scalability and coverage under a different KG schema. We randomly sample 5,000 from the corpus ($\sim$2.2M).
@@
  • QALD-10en [25] (333 test QA pairs) Manually curated, linguistically varied questions over Wikidata; useful for evaluating precision on a small but challenging set. We evaluate on 333 examples.
C.2 Dataset Pre-processing